# The interplay of the parametric instability and MRI in oscillatory shear flows

### v1

@@ 3 Parametric Instability
In order to benchmark the numerical implementation of the setup described in section 2, it is helpful to first compare with a controlled problem for which we have some analytical insight. Therefore, we will start by considering the stability of a low amplitude, free sloshing flow (as described by equations (6), (7) and (11)) which is imposed as an initial condition.
3.1 Summary of the weakly nonlinear PI
Such sloshing flows are known to be unstable due to the parametric instability (PI) (e.g. Gammie et al., 2000), which feeds off the time-dependent, shearing flow. When the shear flow is sufficiently subsonic, with $S\ll 1$, the mechanism can be understood as a three-mode coupling, whereby the sloshing motion acts as the parent mode which destabilizes a pair of inertial waves. The detailed weakly-nonlinear theory of this process is expounded in appendix A, adapting the analysis of Ogilvie & Latter (2013b) and Fairbairn & Ogilvie (2023) to our present setup. Here, we will simply summarize the main results.
Figure 2: Dispersion relation curves tracing the frequency vs radial wavenumber relation for inertial waves supported by our unstratified shearing box. The colored lines denote different branches, labeled by the vertical modal order $n$ (0: blue, 1: orange, 2: green, 3: red, 4: purple, 5: brown). The connective dashed lines indicate the resonant couplings susceptible to parametric instability.
The shearing box supports a variety of wave modes, wherein perturbed variables follow the modal ansatz $X^{\prime}=X_{n}\exp[i(k_{x}x+k_{b}nz-\omega t)]$. When $S=0$, these modes are decoupled and follow a dispersion relation (described by equation (A18)). This is visualized for inertial waves in Fig. 2. The different vertical orders are traced out by the family of colored lines. For example, the orange line denotes $n=1$, which jumps to higher $n$ branches with increasing frequencies, for a given $k_{x}$. However, the introduction of sloshing facilitates a coupling between modes that can lead to instability. The dashed lines indicate mode couplings which resonate effectively with the background sloshing motion. This requires that the inertial waves be of neighboring order in $n$ and are separated in frequency by $\Omega_{0}$. When these conditions are satisfied, energy from the shear flow is efficiently pumped into the inertial waves which grow according to
@@ 3.2 Floquet Theory
The predictions of the weakly-nonlinear theory can be independently verified by appealing to Floquet theory. Indeed, the linearized equations (A12)-(A15), which govern the stability of the sloshing flow, can be written in the form
@@
where $\mathbf{V}$ is a vector of length $4(2N_{\rm max}+1)$ containing the modal, perturbation coefficients $[u_{n},v_{n},w_{n},h_{n}]$, truncated at some order in $|n|\leq N_{\rm max}$. Meanwhile, $\mathsf{N(t)}$ is the matrix encoding the ODE coefficients which are periodic over the dynamical timescale $T=2\pi/\Omega_{0}$. We integrate this equation for each of the unit basis vectors, spanning all possible initial conditions for $\mathbf{V}(0)$. The resulting vectors after one sloshing period, $\mathbf{V}(T)$, can be arranged as the columns of the monodromy matrix $\mathsf{M}(T)$. The eigenvalues of this matrix are the characteristic multipliers $e^{\sigma T}$, whilst the associated eigenvectors $\mathbf{V}_{\rm f}$ define the modal initial conditions which evolve according to each multiplier. Specifically, there exists a solution to equation (19) which evolves according to $\mathbf{V}(t+T)=\mathbf{V}(t)e^{\sigma T}$ where $\mathbf{V}$ is periodic over $T$ and $\mathbf{V}(0)=\mathbf{V}_{\rm f}$.
@@
Let us test this methodology against the predictions of section 3.1. For the applicability of weakly nonlinear theory, we require a small sloshing amplitude and thus adopt $S=0.01$. We then truncate our system of Floquet equations to modes with $|n|\leq 20$ and solve for the monodromy matrix, across a range of $k_{x}$. For each wavenumber, we extract the maximum characteristic multiplier and associated growth rate, $\sigma$. The results are plotted in Fig. 3. The black lines show the results from the Floquet analysis, which exhibit a series of unstable bands, which coincide with the resonant wavelengths predicted by the dashed lines in Fig. 2. Meanwhile, the red markers plot the growth rates predicted from equation (18) of the three-mode coupling theory. Clearly there is very good agreement in this weak sloshing regime, mutually verifying our analysis and the Floquet implementation. We have also tested (not shown here) that the Floquet growth rates vary linearly in $S$, in accordance with theory, provided that the flows remain in the weak sloshing regime. Nonetheless, it should be emphasized that the applicability of Floquet theory extends beyond this weakly nonlinear regime, since the problem structure follows equation (19) for arbitrary values of $S$. In fact, this will provide a useful tool to compare with our simulations in sections 4 and 5, where the sloshing magnitude can become of sonic order. However, at such large $S$ one should be careful to include sufficient modes in the Floquet analysis to ensure convergence. Indeed, the strong sloshing facilitates enhanced modal communication, which will spread the couplings beyond three-waves. Furthermore, there is the possibility that there are no modal solutions for very large shear flows, as discussed by Ogilvie & Latter (2013b) and Paardekooper & Ogilvie (2019), which would preclude convergence entirely.
@@ 3.3 Numerical verification
We choose our box size judiciously so that it permits the unstable radial wavenumbers predicted from the three-mode coupling theory. In particular, we will test the growth of the longest radial wavelength resonance, as shown by the dashed orange line in Fig. A18, which occurs at $k_{x}H=15.6$ and connects the $n=1$ branch at $\omega/\Omega_{0}=-0.373$, with the $n=2$ branch at $\omega/\Omega_{0}=0.627$. In order to allow three radial wavelengths to fit in the box, we set $L_{x}=1.2084$. Meanwhile, we choose $L_{y}=\pi$ and $L_{z}=H=1$. Capturing the growth rate accurately requires a sufficient resolution, so we establish a grid with $(N_{x},N_{y},N_{z})=(154,200,128)$ which well resolves the vertical and radial wavelengths. We then seed the instability by adding the resonant pair of inertial modes, each with eigenmode structure $\mathbf{V}_{n}^{(0)}(\omega;k_{x}H)$ described by equation (A19). Each mode is normalized so that the maximal radial velocity associated with it is 0.0001 – a factor of 100 smaller than the sloshing. We then run the simulation on 1 H200 GPU for 60 orbital periods.
@@
Figure 4: Numerical simulation verifying three-mode coupling parametric instability. Upper panel: the initialized modal structure in $u_{z}^{\prime}$ is shown as the heat map, over-plotted with white arrows denoting the meridional streamlines. Lower panel: the growth of the vertical kinetic energy in our simulation (black line) is fit (red dashed line) between $0<t<30T$ in order to extract the energy growth rate $0.0189652\sigma=$ 0.018965.
@@
As desired, our numerical experiment nicely captures the modal amplification by parametric instability. The key results are visualized in Fig. 4. The heat map in the upper panel shows the initialized modal structure of the vertical velocity perturbation $u_{z}^{\prime}$. Meanwhile, the meridional streamlines are over-plotted as the white arrows. These oblique vertical flow structures are also typical signatures of the PI which manifest in other disk setups (e.g. Ogilvie & Latter, 2013b; Deng et al., 2020). In the lower panel, the black line plots the logarithm of the total vertical component of the kinetic energy $E_{{\rm k},z}$ summed over the box. This undergoes a clear exponential rise phase which is well fit by the red dashed line between $0<t<30T$. The slope of this line indicates an energy growth rate of $0.018965$ which yields an amplitude growth rate, $\sigma=0.0094825$. Pleasingly, this is very close to the predictions of the three-mode coupling theory which gives $\sigma_{\rm WN}=0.00978919$ (as plotted by the first red marker in Fig. 3). One might rush to concluding that the slight discrepancy here is due to numerical dissipation. However, we have also performed tests with double and quadrupole the grid resolution and confirm that our fiducial setup is already suitably converged. Indeed, if we initialize a free wave and switch off the free sloshing, we can track its numerical dissipation which is found to have an energy damping rate of $\sim 0.0001<|\sigma_{\rm WN}-\sigma|$. This is too small to explain the aforementioned difference between the predicted and measured growth rates. Instead, the slight discrepancy is mainly attributed to subdominant, direct nonlinear interactions between the two waves, which are neglected in the weakly nonlinear analysis. Indeed, the smooth turnover in $\ln(E_{\rm k,z})$ around $t=50T$ is also indicative of these nonlinear, secular effects. Additional tests, using waves initialized with smaller seed amplitudes, show this variation over longer timescales and the growth rates converge towards the weakly nonlinear predictions.
@@ 4 Hydrodynamic results
Having outlined the theoretical setup in section 2, and benchmarked the numerics with a controlled problem in section 3, we are now in a position to examine the role of hydrodynamic instability, and the ensuing turbulence, in setting an equilibrium state when the sloshing is forced.
@@ 4.1 Numerical setup
The numerical setup is similar to that described in section 3.3 but now we forgo the free sloshing initial condition (setting $S=0$) and instead employ the forcing function described by equation (5) as a user-implemented source term in AthenaK. Our fiducial study will adopt a value of $\psi_{\rm f}=0.3$ which is large enough to set up a clear sloshing signal atop the turbulence. We will compare this with smaller forcing values in section 5.4. Our box size is taken to be $(L_{x},L_{y},L_{z})=(9.6672,\pi,1)$ which is elongated by a factor of 8 in the $x$ dimension compared to the experiment performed in section 3.3. As we will show in the following results, the forced sloshing, turbulent state wants to select a longer radial length scale compared with the weakly nonlinear theory. Indeed, our early tests found that the diagnostics for the equilibrium sloshing state were sensitive to the box size, before converging for sufficiently wide boxes. To maintain the same resolution as before, we also increase the number of cells so that $(N_{x},N_{y},N_{z})=(1232,200,128)$. We start off from a Keplerian flow, with white-noise fluctuations in the density and velocity components at the level of 0.01, which act to seed the PI. We then run the experiment for $60T$ which tracks the initial resonant growth phase, the disruption due to the PI and the subsequent equilibration toward a steady sloshing state moderated by turbulent hydrodynamic stresses. Throughout the analysis of our simulations, we will regularly take averages of quantities across our computational domain. We will denote spatial averages of a quantity $X$ as
[…]
@@ 4.2 Growth of instability
Figure 5: Growth of the PI for our $\psi_{\rm f}=0.3$ hydro run. Panel (a): Epicylic kinetic energy measured in the simulation (black line) and the result expected for the inviscid, laminar response (red dashed line). Inset panel: the resonant growth of the sloshing motion $A_{\rm res}$ during early times. Panel (b): the logarithm of the box-integrated vertical kinetic energy.
In Fig. 5 we track the evolution of the forced sloshing motion as a function of time. Specifically, in panel (a) we are plotting the box-integrated epicylic kinetic energy which is defined as
@@
where $C=A^{2}+4B^{2}$ is computed from equations (13)-(14). For example, the inset panel shows the growing response of the $x$ oscillation $A_{\rm res}$ over the first two orbital periods. At early times, we see very good agreement between these two curves as the epicylic motion is amplified in accordance with the laminar, resonant theory. However, at $t/T\sim 2$ we see the simulation result turnover as the epicylic amplitude is strongly quenched. The PI efficiently siphons energy from the coherent sloshing motion and redirects it towards growing inertial modes. This is clearly seen in panel (b) which shows the logarithm of the box-integrated vertical kinetic energy. This undergoes an approximately exponential growth phase, before saturating just after the peak in epicylic energy. At late times, both $E_{\rm epi}$ and $E_{{\rm k},z}$ settle down to a quasi-steady state. This represents the equilibrium sloshing state in which the coherent epicylic motions have come into balance with the nonlinear manifestation of the PI.
@@
Figure 6: PI excited modal structure. Panel (a): meridional $x-z$ cross-section of $\{u_{z}\}_{y}$. Panel (b): maximal Floquet growth rates extracted as a function of the $k_{x}$ (at the allowed values quantized by the box) for different sloshing amplitudes $S=0.5$ (blue), $1.0$ (black) and $1.5$ (red). Panel (c): the modal structure corresponding to the maximal growth rate for $S=1.0$. The $u_{n}$, $v_{n}$, $w_{n}$ and $h_{n}$ contributions are shown as blue circles, orange squares, green triangles and red crosses respectively.
@@
One should notice from the inset panel in Fig. 5, that the radial flows become supersonic around $t\sim T$. This strong, and time-dependent, sloshing amplitude obviously departs the realms of applicability for the three-mode coupling analysis discussed in section 3.1. Therefore, we might expect the emergent PI to exhibit different properties. Indeed, the exponential growth rate of the vertical velocity in panel (b) is not as clean as that found in the controlled experiment shown in Fig. 4. Indeed, as $S$ becomes large one should expect a transient modulation of the growing waves over a sloshing cycle. Moreover, the spatial structure of the growing perturbations is notably different. This is visualized in Fig. 6 where panel (a) plots the density-weighted, $y$-averaged vertical velocity perturbations $\{u_{z}\}_{y}$ across the meridional extent of the box at $t/T=1.5$. This shows a striking pattern of alternating vertical velocity channels that have a predominantly $n=0$ vertical structure, reminiscent of that found in Fig. 6 of Fairbairn & Ogilvie (2023). Hereafter we will refer to these distinctive structures as ‘elevator’ flows. [^3] Similar upwards and downwards velocity channels have curiously emerged in a variety of other unstratified local simulation setups (e.g. Dewberry et al., 2020; Teed & Latter, 2021), wherein they are vaguely attributed to generic forcing by inertial wave turbulence. In this snapshot, we also see that the pattern fits around 6 radial wavelengths into the box, which corresponds to a wavenumber of $k_{x}H=3.9$. This is clearly a much longer wavelength compared with the smallest resonant wavenumber seen in the weakly nonlinear theory, exemplified in Fig. 3. This once again underlines the importance of simulating a wide enough box that captures a number of these wavelengths. Upon testing different $L_{x}$, we find that the emergence of a preferred wavelength (which can change with time) is a robust result; supporting a physical origin over some scaling with the box width.
@@
To better understand the emergence of the elevator flows, we can appeal to the use of the Floquet theory (see section 3.2), which can still be applied for large values of $S$. Whilst this formalism strictly requires a time-periodic set of equations for which $S$ is constant, we find that it still maps effectively onto our full simulation results. In panel (b) of Fig. 6 we plot the maximum growth rate computed from the Floquet theory as a function of $k_{x}$ using $|n|\leq 30$ modes. Specifically, we extract these growth rates at the discrete values admitted by our periodic box. Since the sloshing amplitude increases with time during the resonant growth phase, we perform this analysis for three different values of $S=0.5$, $1.0$ and $1.5$ (plotted by the blue, black and red markers respectively). We observe that the narrow resonant peaks, found in the three-mode coupling theory, are replaced by extended bands of instability (also see Ogilvie & Latter, 2013b; Paardekooper & Ogilvie, 2019). Moreover, we find peaks at much smaller wavenumbers, as indicated by the accordingly colored, dashed vertical lines. As the sloshing increases, the optimal mode moves towards longer wavelengths. This matches the behavior seen in the simulation where we observe small-scale structure at early times, before longer wavelengths begin to dominate later. The optimal wavelength for $S=1.0$ occurs at $k_{x}H=3.9$ which is in agreement with the snapshot in panel (a). Although the radial sloshing first reaches $A_{\rm res}=1$ around $t/T\sim 1$, whilst panel (a) actually corresponds to $t/T=1.5$, it is reasonable to expect that there is some time delay for the preferred mode to grow to prominence. Of course, since the sloshing amplitude is continually changing, the preferred mode also evolves. Eventually during the quasi-steady sloshing state the vertical velocity settles into a structure with 4 radial wavelengths in the box. Finally, in panel (c) we plot the eigenmode structure for the $S=1.0$ optimal mode. The contribution from each of the vertical expansion coefficients is plotted as a function of $n$. The $w_{0}$ mode dominates the structure as the green, triangle markers peak at $n=0$. This pleasingly supports the development of the prominent $n=0$ elevator flows found in the simulations. These vertical flows enhance the communication between the horizontal, sloshing layers – leading to hydrodynamic stresses which provide a rigidity, resisting the resonant forcing (see sections 4.3 and 4.4).

### v2

@@ 3 Parametric Instability
In order to benchmark the numerical implementation of the setup described in Section 2, it is helpful to first compare with a controlled problem for which we have some analytical insight. Therefore, we will start by considering the stability of a low-amplitude, inviscid, free sloshing flow (as described by equations (6), (7) and (11)), which is imposed as an initial condition.
3.1 Summary of the Weakly Nonlinear Parametric Instability
Such sloshing flows are known to be unstable due to the parametric instability (PI) (e.g. Gammie et al., 2000), which feeds off the time-dependent, shearing flow. When the shear flow is sufficiently subsonic, with $S\ll 1$, the mechanism can be understood as a three-mode coupling, whereby the sloshing motion acts as the parent mode which destabilizes a pair of inertial waves. The detailed weakly nonlinear theory of this process is expounded in appendix A, adapting the analysis of Ogilvie & Latter (2013b) and Fairbairn & Ogilvie (2023) to our present setup. Here, we will simply summarize the main results.
Figure 2: Dispersion relation curves tracing the frequency vs radial wavenumber relation for locally axisymmetric inertial waves supported by our unstratified shearing box. The colored lines denote different branches, labeled by the vertical modal order $n$ (0: blue, 1: orange, 2: green, 3: red, 4: purple, 5: brown). The connective dashed lines indicate the resonant couplings susceptible to parametric instability.
The shearing box supports a variety of wave modes, wherein perturbed variables follow the modal ansatz $X^{\prime}=X_{n}\exp[i(k_{x}x+k_{\rm b}nz-\omega t)]$. When $S=0$, these modes are decoupled and follow a dispersion relation (described by equation (A18)). This is visualized for inertial waves in Fig. 2. The different vertical orders are traced out by the family of colored lines. For example, the orange line denotes $n=1$, which jumps to higher $n$ branches with increasing frequencies, for a given $k_{x}$. However, the introduction of sloshing facilitates a coupling between modes that can lead to instability. The dashed lines indicate mode couplings that resonate effectively with the background sloshing motion. This requires that the inertial waves be of neighboring order in $n$ and are separated in frequency by $\Omega_{0}$, the sloshing frequency as seen in the coorbital frame. When these conditions are satisfied, energy from the shear flow is efficiently pumped into the inertial waves, which grow according to
@@ 3.2 Floquet Theory
The predictions of the weakly nonlinear theory can be independently verified by appealing to Floquet theory. Indeed, the linearized equations (A12)-(A15), which govern the stability of the sloshing flow, can be written in the form
@@
where $\mathbf{V}$ is a vector of length $4(2N_{\rm max}+1)$ containing the modal, perturbation coefficients $u_{n},v_{n},w_{n},h_{n}]$, truncated at some order in $|n|\leq N_{\rm max}$. Meanwhile, $\mathsf{N(t)}$ is the matrix encoding the ordinary differential equation coefficients, which are periodic over the dynamical timescale $T=2\pi/\Omega_{0}$. We integrate this equation for each of the unit basis vectors, spanning all possible initial conditions for $\mathbf{V}(0)$. The resulting vectors after one sloshing period, $\mathbf{V}(T)$, can be arranged as the columns of the monodromy matrix $\mathsf{M}(T)$. The eigenvalues of this matrix are the characteristic multipliers $e^{\sigma T}$, whilst the associated eigenvectors $\mathbf{V}_{\rm f}$ define the modal initial conditions which evolve according to each multiplier. In particular, there exists a solution to equation (19), which evolves according to $\mathbf{V}(t+T)=\mathbf{V}(t)e^{\sigma T}$ where $\mathbf{V}$ is periodic over $T$ and $\mathbf{V}(0)=\mathbf{V}_{\rm f}$.
@@
Let us test this methodology against the predictions of Section 3.1. For the applicability of weakly nonlinear theory, we require a small sloshing amplitude and thus adopt $S=0.01$. We then truncate our system of Floquet equations to modes with $|n|\leq 20$ and solve for the monodromy matrix across a range of $k_{x}$. For each wavenumber, we extract the maximum characteristic multiplier and associated growth rate, $\sigma$. The results are plotted in Fig. 3. The black lines show the results from the Floquet analysis, which exhibit a series of unstable bands that coincide with the resonant wavelengths predicted by the dashed lines in Fig. 2. Meanwhile, the red markers plot the growth rates predicted from equation (18) of the three-mode coupling theory. Clearly, there is very good agreement in this weak sloshing regime, mutually verifying our analysis and the Floquet implementation. We have also tested (not shown here) that the Floquet growth rates vary linearly in $S$, in accordance with theory, provided that the flows remain in the weak sloshing regime. Nonetheless, it should be emphasized that the applicability of Floquet theory extends beyond this weakly nonlinear regime, since the problem structure follows equation (19) for arbitrary values of $S$. In fact, this will provide a useful tool to compare with our simulations in sections 4 and 5, where the sloshing magnitude can become of sonic order. However, at such a large $S$, one should be careful to include sufficient modes in the Floquet analysis to ensure convergence. Indeed, the strong sloshing facilitates enhanced modal communication, which will spread the couplings beyond three waves. Furthermore, there is the possibility that there are no modal solutions for very large shear flows, as discussed by Ogilvie & Latter (2013b) and Paardekooper & Ogilvie (2019), which would preclude convergence entirely.
@@ 3.3 Numerical Verification
We choose our box size judiciously so that it permits the unstable radial wavenumbers predicted from the three-mode coupling theory. In particular, we will test the growth of the longest radial wavelength resonance, as shown by the dashed orange line in Fig. 2, which occurs at $k_{x}H=15.6$ and connects the $n=1$ branch at $\omega/\Omega_{0}=-0.373$, with the $n=2$ branch at $\omega/\Omega_{0}=0.627$. In order to allow for three radial wavelengths to fit in the box, we set $L_{x}=1.2084$. Meanwhile, we choose $L_{y}=\pi$ and $L_{z}=H=1$. Capturing the growth rate accurately requires a sufficient resolution, so we establish a grid with $(N_{x},N_{y},N_{z})=(154,200,128)$ which well resolves the vertical and radial wavelengths. We then seed the instability by adding the resonant pair of inertial modes, each with eigenmode structure $\mathbf{V}_{n}^{(0)}(\omega;k_{x}H)$ described by equation (A19). Each mode is normalized so that the maximal radial velocity associated with it is 0.0001 – a factor of 100 smaller than the sloshing. We then run the simulation on 1 H200 GPU for 60 orbital periods.
@@
Figure 4: Numerical simulation verifying three-mode coupling parametric instability. Upper panel: The initialized modal structure in $u_{z}^{\prime}$ is shown as the heat map, overplotted with white arrows denoting the meridional streamlines. Lower panel: The growth of the vertical kinetic energy in our simulation (black line) is fit (red dashed line) between $0<t<30T$ in order to extract the energy growth rate $2\sigma=0.018965$.
@@
As desired, our numerical experiment nicely captures the modal amplification by parametric instability. The key results are visualized in Fig. 4. The heat map in the upper panel shows the initialized modal structure of the vertical velocity perturbation $u_{z}^{\prime}$. Meanwhile, the meridional streamlines are overplotted as the white arrows. These oblique vertical flow structures are also typical signatures of the PI, which manifest in other disk setups (e.g. Ogilvie & Latter, 2013b; Deng et al., 2020). In the lower panel, the black line plots the logarithm of the total vertical component of the kinetic energy $E_{{\rm k},z}$ summed over the box. This undergoes a clear exponential rise phase, which is well fit by the red dashed line between $0<t<30T$. The slope of this line indicates an energy growth rate of $0.018965$, which yields an amplitude growth rate of $\sigma=0.0094825$. Pleasingly, this is very close to the predictions of the three-mode coupling theory, which gives $\sigma_{\rm WN}=0.00978919$ (as plotted by the first red marker in Fig. 3). One might rush to a conclusion that the slight discrepancy here is due to numerical dissipation. However, we have also performed tests with double and quadrupole the grid resolution and confirm that our fiducial setup is already suitably converged. Indeed, if we initialize a free wave and switch off the free sloshing, we can track its numerical dissipation, which is found to have an energy damping rate of $\sim 0.0001<|\sigma_{\rm WN}-\sigma|$. This is too small to explain the aforementioned difference between the predicted and measured growth rates. Instead, the slight discrepancy is mainly attributed to subdominant, direct nonlinear interactions between the two waves, which are neglected in the weakly nonlinear analysis. Indeed, the smooth turnover in $\ln(E_{\rm k,z})$ around $t=50T$ is also indicative of these nonlinear, secular effects. Additional tests, using waves initialized with smaller seed amplitudes, show this variation over longer timescales, and the growth rates converge toward the weakly nonlinear predictions.
@@ 4 Hydrodynamic results
Having outlined the theoretical setup in Section 2, and benchmarked the numerics with a controlled problem in Section 3, we are now in a position to examine the role of hydrodynamic instability and the ensuing turbulence in setting an equilibrium state when the sloshing is forced.
@@ 4.1 Numerical Setup
The numerical setup is similar to that described in Section 3.3, but now we forgo the free sloshing initial condition (setting $S=0$) and instead employ the forcing function described by equation (5) as a user-implemented source term in AthenaK. Our fiducial study will adopt a value of $\psi_{\rm f}=0.3$ which is large enough to set up a clear sloshing signal atop the turbulence. We will compare this with smaller forcing values in Section 5.4. Once again, we set the explicit viscosity to be 0 for these hydrodynamic runs. In appendix B we also perform a number of viscous experiments, for comparison with our MHD runs (see Section 5) which necessarily include nonideal dissipation in order to obtain a converged MRI. Our box size is taken to be $(L_{x},L_{y},L_{z})=(9.6672,\pi,1)$, which is elongated by a factor of 8 in the $x$ dimension compared to the experiment performed in Section 3.3. As we will show in the following results, the forced sloshing, turbulent state wants to select a longer radial length scale compared with the weakly nonlinear theory. Indeed, our early tests found that the diagnostics for the equilibrium sloshing state were sensitive to the box size, before converging for sufficiently wide boxes. This also allows for a finer sampling of quantized radial wavenumber, hence minimizing the possibility that certain mode couplings are missed. To maintain the same resolution as before, we also increase the number of cells so that $(N_{x},N_{y},N_{z})=(1232,200,128)$. We start off from a Keplerian flow, with white-noise fluctuations in the density and velocity components at the level of 0.01, which act to seed the PI. We then run the experiment for $60T$, which tracks the initial resonant growth phase, the disruption due to the PI, and the subsequent equilibration toward a steady sloshing state moderated by turbulent hydrodynamic stresses. Throughout the analysis of our simulations, we will regularly take averages of quantities across our computational domain. We will denote spatial averages of a quantity $X$ as
$\langle{X\rangle}_{\Pi_{i}x_{i}}=\frac{\int X\,dD}{\int dD},$ (20)
@@ 4.2 Growth of Instability
Figure 5: Growth of the PI for our $\psi_{\rm f}=0.3$ hydrodynamic run. Panel (a): epicylic kinetic energy measured in the simulation (black line) and the result expected for the inviscid, laminar response (red dashed line). Inset panel: the resonant growth of the sloshing motion $A_{\rm res}$ during early times. Panel (b): the logarithm of the box-integrated vertical kinetic energy.
In Fig. 5 we track the evolution of the forced sloshing motion as a function of time. In particular, in panel (a), we are plotting the box-integrated epicylic kinetic energy, which is defined as
@@
where $C=A^{2}+4B^{2}$ is computed from equations (13)-(14). For example, the inset panel shows the growing response of the $x$ oscillation $A_{\rm res}$ over the first two orbital periods. At early times, we see very good agreement between these two curves as the epicylic motion is amplified in accordance with the laminar, resonant theory. However, at $t/T\sim 2$ we see the simulation result turnover as the epicylic amplitude is strongly quenched. The PI efficiently siphons energy from the coherent sloshing motion and redirects it toward growing inertial modes. This is clearly seen in panel (b), which shows the logarithm of the box-integrated vertical kinetic energy. This undergoes an approximately exponential growth phase, before saturating just after the peak in epicylic energy. At late times, both $E_{\rm epi}$ and $E_{{\rm k},z}$ settle down to a quasi-steady state. This represents the equilibrium sloshing state in which the coherent epicylic motions have come into balance with the nonlinear manifestation of the PI.
@@
Figure 6: PI excited modal structure. Panel (a): meridional xxzz cross-section of $\{u_{z}\}_{y}$. Panel (b): maximal Floquet growth rates extracted as a function of the $k_{x}$ (at the allowed values quantized by the box) for different sloshing amplitudes $S=0.5$ (blue), $1.0$ (black) and $1.5$ (red). Panel (c): the modal structure corresponding to the maximal growth rate for $S=1.0$. The $u_{n}$, $v_{n}$, $w_{n}$ and $h_{n}$ contributions are shown as blue circles, orange squares, green triangles, and red crosses, respectively.
@@
One should notice from the inset panel in Fig. 5, that the radial flows become supersonic around $t\sim T$. This strong, and time-dependent, sloshing amplitude obviously departs from the realms of applicability for the three-mode coupling analysis discussed in Section 3.1. Therefore, we might expect the emergent PI to exhibit different properties. Indeed, the exponential growth rate of the vertical velocity in panel (b) is not as clean as that found in the controlled experiment shown in Fig. 4. Indeed, as $S$ becomes large, one should expect a transient modulation of the growing waves over a sloshing cycle. Moreover, the spatial structure of the growing perturbations is notably different. This is visualized in Fig. 6, where panel (a) plots the density-weighted, $y$-averaged vertical velocity perturbations $\{u_{z}\}_{y}$ across the meridional extent of the box at $t/T=1.5$. This shows a striking pattern of alternating vertical velocity channels that have a predominantly $n=0$ vertical structure, reminiscent of that found in Fig. 6 of Fairbairn & Ogilvie (2023). Hereafter, we will refer to these distinctive structures as ‘elevator’ flows. In this snapshot, we also see that the pattern fits around 6 radial wavelengths into the box, which corresponds to a wavenumber of $k_{x}H=3.9$. This is clearly a much longer wavelength compared with the smallest resonant wavenumber seen in the weakly nonlinear theory, exemplified in Fig. 3. This once again underlines the importance of simulating a wide enough box that captures a number of these wavelengths. Upon testing different $L_{x}$, we find that the emergence of a preferred wavelength (which can change with time) is a robust result, supporting a physical origin over some scaling with the box width.
@@
To better understand the emergence of the elevator flows, we can appeal to the use of the Floquet theory (see Section 3.2), which can still be applied for large values of $S$. Whilst this formalism strictly requires a time-periodic set of equations for which $S$ is constant, we find that it still maps effectively onto our full simulation results. In panel (b) of Fig. 6, we plot the maximum growth rate computed from the Floquet theory as a function of $k_{x}$ using $|n|\leq 30$ modes. In particular, we extract these growth rates at the discrete values admitted by our periodic box. Since the sloshing amplitude increases with time during the resonant growth phase, we perform this analysis for three different values of $S=0.5$, $1.0$, and $1.5$ (plotted by the blue, black, and red markers, respectively). We observe that the narrow resonant peaks, found in the three-mode coupling theory, are replaced by extended bands of instability (also see Ogilvie & Latter, 2013b; Paardekooper & Ogilvie, 2019). Moreover, we find peaks at much smaller wavenumbers, as indicated by the accordingly colored, dashed vertical lines. As the sloshing increases, the optimal mode moves toward longer wavelengths. This matches the behavior seen in the simulation, where we observe small-scale structure at early times, before longer wavelengths begin to dominate later. The optimal wavelength for $S=1.0$ occurs at $k_{x}H=3.9$, which is in agreement with the snapshot in panel (a). Although the radial sloshing first reaches $A_{\rm res}=1$ around $t/T\sim 1$, whilst panel (a) actually corresponds to $t/T=1.5$, it is reasonable to expect that there is some time delay for the preferred mode to grow to prominence. Of course, since the sloshing amplitude is continually changing, the preferred mode also evolves. Eventually, during the quasi-steady sloshing state, the vertical velocity settles into a structure with four radial wavelengths in the box. Finally, in panel (c), we plot the eigenmode structure for the $S=1.0$ optimal mode. The contribution from each of the vertical expansion coefficients is plotted as a function of $n$. The $w_{0}$ mode dominates the structure as the green, triangle markers peak at $n=0$. This pleasingly supports the development of the prominent $n=0$ elevator flows found in the simulations. These vertical flows enhance the communication between the horizontal, sloshing layers – leading to hydrodynamic stresses which provide a rigidity, resisting the resonant forcing (see sections 4.3 and 4.4).
Similar upward and downward velocity channels have curiously emerged in a variety of other unstratified local simulation setups (e.g. Dewberry et al., 2020; Teed & Latter, 2021), wherein they are vaguely attributed to generic forcing by inertial wave turbulence. These elevator modes might also be tentatively connected with the spectrum for oceanic internal waves, which is found to peak around the lowest frequency modes (Garrett & Munk, 1979). [^3] We thank the referee for bringing this connection to our attention. The physical picture is based on the preponderance for resonant triad interactions, which transfer energy from parent modes to children with lower frequencies (Müller et al., 1986; Shavit et al., 2025). This weakly nonlinear wave turbulence can therefore establish an inverse cascade toward low-frequency modes. Whether this speculative mechanism helps to sustain the elevator modes in our context is beyond the scope of our present study, but ripe for future investigation.